40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 14040     Entry time: Thu Jul 5 17:58:04 2018
Author: keerthana, sandrine 
Type: Update 
Category: elog 
Subject:  

(Analisa, Sandrine, Keerthana)

Today Annalisa helped us to understand the new set up used to make the frequency scans of the AUX laser. While tracking the cables it seemed that there were quite a lot of cables near the mixer. So we have reconnected one of the splitter which was splitting the RF out put signal from the Agilent and have placed it just near the Agilent itself. A picture of the changed setup is provided below. The splitter divides the signal into two components. One goes to the LO port of the mixer and the other goes to the R port of the Agilent. We have tried locking the PLL after the change and it works fine. We are trying to make a diagram of the setup now, which we will upload shortly.

 

Attachment 1: setup1.jpg  3.586 MB  Uploaded Thu Jul 5 19:18:30 2018  | Hide | Hide all
setup1.jpg
Attachment 2: setup2.jpg  3.216 MB  Uploaded Thu Jul 5 19:18:36 2018  | Hide | Hide all
setup2.jpg
ELOG V3.1.3-